Can someone help me on this?
The length of a rectangular garden is 2 ft. less than 3 times its width. If the perimeter of the garden is 44 ft, find the dimensions of the garden.
Thanks, Lyn

Answer by checkley71(8403)

LENGTH=3W-2
WIDTH=W
PERIMETER=2*LENGTH+2*WIDTH
44=2(3W-2)+2W
44=6W-4+2W
44+4=8W
48=8W
W=48/8
W=6 & L=3*6-2 OR L=18-2 L=16
PROOF
44=2*16+2*6
44=32+12
44=44